OpenStudy (anonymous):

can i have someone help me understand and show me how to do a problem like this? 1/3+(-1/2)+5/6

hero (hero):

Now find the least common multiple between 2,3 and 6

OpenStudy (anonymous):

im sorry the 1 is a 2

hero (hero):

Which one is a two?

OpenStudy (anonymous):

1/3+(-2/3)+5/6

hero (hero):

\[\frac{1}{3} - \frac{2}{3} + \frac{5}{6}\]

OpenStudy (anonymous):

yea

hero (hero):

1. What is the least common multiple between 3 and 6? Do you know it?

OpenStudy (anonymous):

6

hero (hero):

Very good

hero (hero):

So now we have to make the denominators of the other two fractions six as well by creating equivalent fractions.

OpenStudy (anonymous):

so all the bottom is 6 correct

hero (hero):

And replacing those fractions with the equivalent ones. To do that set up a proportion like these: \[\frac{1}{3} = \frac{x}{6}\]\[\frac{2}{3} = \frac{y}{6}\] Now find x and y

hero (hero):

Yes, eventually all of the bottoms (denominators) of the fractions will be six.

OpenStudy (anonymous):

ok

hero (hero):

Can you find x and y for me please?

OpenStudy (anonymous):

i have never seen it set up like you just showed me

hero (hero):

What I wrote above are proportions. What grade are you in?

OpenStudy (anonymous):

college

hero (hero):

You should know what a proportion is by now and you should know how to solve them.

hero (hero):

For both of them all you do is cross multiply to get: 6 = 3x and 12 = 3y

OpenStudy (anonymous):

should but really never seen it that way

hero (hero):

It's just a way to set up and solve equivalent fractions

OpenStudy (anonymous):

i think my teacher showed us another way

hero (hero):

Probably, but it's really not much different from what I am showing.

OpenStudy (anonymous):

so it would be 6*1 and 6*2

hero (hero):

Maybe someone else can help you. I don't think you're getting this.

OpenStudy (anonymous):

oh well whatever i will figure it out